# Self-Supervised Contrastive Learning is Approximately Supervised Contrastive Learning

### v1

@@ 1 Introduction
Unsupervised representation learning refers to a class of algorithmic approaches designed to discover meaningful representations of complex data without relying on explicit supervision signals. The goal is to learn representations that preserve and expose semantic information, allowing them to be effectively leveraged in downstream supervised tasks. In recent years, these methods have proven to be effective in pre-training models on unlabeled data, enabling strong generalization on downstream computer vision tasks (hjelm2018learning, ; oord2019representationlearningcontrastivepredictive, ; NEURIPS2019_ddf35421, ; NEURIPS2020_70feb62b, ; pmlr-v119-chen20j, ; chen2020improvedbaselinesmomentumcontrastive, ; Chen_2021_ICCV, ; hénaff2020dataefficientimagerecognitioncontrastive, ; li2021prototypical, ; misra2019selfsupervisedlearningpretextinvariantrepresentations, ; 10.1007/978-3-030-58621-8_45, ; NEURIPS2020_4c2e5eaa, ; zbontar2021barlow, ; bardes2022vicreg, ; he2022masked, ; oquab2024dinov2learningrobustvisual, ) and in natural language processing (fang2020certcontrastiveselfsupervisedlearning, ; NEURIPS2020_1457c0d6, ; gao-etal-2021-simcse, ; su-etal-2022-tacl, ; reimers-gurevych-2019-sentence, ; zhang-etal-2021-pairwise, ; giorgi-etal-2021-declutr, ; yan-etal-2021-consert, ).
@@
Figure 1: DCL forms semantic clusters without label supervision, while NSCL yields tighter, more separable clusters, despite not explicitly pulling same-class samples together. We plot UMAP visualizations for (top) decoupled contrastive learning (DCL) (10.1007/978-3-031-19809-0_38, ) and (bottom) negatives-only supervised contrastive learning (NSCL) training on mini-ImageNet.
One of the most successful paradigms for unsupervised learning is self-supervised contrastive learning (CL), where models are trained to distinguish different augmented views of the same image (positives) from views of other images (negatives), typically by minimizing the InfoNCE loss (pmlr-v9-gutmann10a, ; oord2019representationlearningcontrastivepredictive, ). Notable methods in this category include SimCLR (pmlr-v119-chen20j, ), MoCo (He_2020_CVPR, ; chen2020improvedbaselinesmomentumcontrastive, ; Chen_2021_ICCV, ), and CPC (oord2019representationlearningcontrastivepredictive, ). These approaches have led to representations that generalize remarkably well to downstream tasks, often competing with or even surpassing supervised learning. For example, an ImageNet-1M pre-trained MoCo achieves 81.5 $\mathrm{AP}_{50}$ on PASCAL VOC (journals/ijcv/EveringhamGWWZ10, ), while its supervised counterpart achieves 81.3 $\mathrm{AP}_{50}$.
Although the representations learned in supervised classification problems are relatively well understood–thanks to characterizations such as neural collapse doi:10.1073/pnas.2015509117 ; han2022neural , which occurs under a variety of loss functions including cross-entropy (LU2022224, ; 10.5555/3692070.3692468, ; pmlr-v139-graf21a, ), mean squared error (MSE) (han2022neural, ; tirer2022extendedunconstrainedfeaturesmodel, ; zhou2022optimizationlandscapeneuralcollapse, ), and supervised contrastive loss (pmlr-v139-graf21a, ; pmlr-v162-awasthi22b, ; gill2023engineeringneuralcollapsegeometry, ; kini2024symmetric, )–the theoretical understanding of SSL remains limited. Surprisingly, despite the absence of labels, SSL-trained models often produce representations that closely resemble those learned via supervised training: they tend to cluster nicely with respect to semantic classes  (amir2021deep, ; shaul2023reverse, ; weng2025clusteringpropertiesselfsupervisedlearning, ) (see also Fig. 1) and support downstream tasks such as linear classification and clustering (pmlr-v119-chen20j, ; He_2020_CVPR, ; chen2020improvedbaselinesmomentumcontrastive, ; Chen_2021_ICCV, ; oord2019representationlearningcontrastivepredictive, ). This raises the following question:
@@
  • Duality between self-supervised and supervised CL: In Thm. 1 we formally establish a connection between decoupled contrastive learning (DCL) (10.1007/978-3-031-19809-0_38, ) and a supervised variant we call negatives-only supervised contrastive loss (NSCL), in which negative pairs are drawn from different classes. Our main insight is that, although DCL does not explicitly exclude same-class samples from the loss denominator, the probability of treating them as negatives vanishes as the number of classes increases. We prove that the gap between the two losses shrinks with more classes. Unlike prior analyses (arora2019theoreticalanalysiscontrastiveunsupervised, ; pmlr-v162-saunshi22a, ; tosh2021contrastive, ; pmlr-v162-awasthi22b, ; pmlr-v195-parulekar23a, ), our result holds without assumptions on model architecture, data distribution, or augmentations. Figs. 23 show that the two losses are highly correlated, their gap decreases as the number of classes increases, and minimizing the DCL loss implicitly minimizes the NSCL loss.
  • Estimating the few-shot error on downstream tasks: Prop. 1 introduces a bound on the $m$-shot classification error based on geometric properties of the learned representation. Specifically, the bound implies that lower class-distance-normalized-variance (CDNV) (see (galanti2022on, ) and Sec. 3.2) and directional CDNV (see Sec. 3.2) lead to lower $m$-shot error, with the impact of the regular CDNV diminishing as the number $m$ of per-class labeled examples increases. Empirically, we verify this behavior for DCL-trained models (Fig. 5) and find the bound to be tight in practice (Fig. 4).
  • Characterizing the global minima of the NSCL loss: In Thm. 2, we show that any global minimizer of the NSCL loss satisfies: (i) Augmentation collapse—all augmented views of a sample map to the same point; (ii) Within-class collapse—all samples from the same class share a representation; and (iii) Simplex ETF structure—class centers form a maximally separated, symmetric configuration. These solutions coincide with the optimal solutions of certain supervised classification settings, including those using cross-entropy loss (LU2022224, ; 10.5555/3692070.3692468, ; pmlr-v139-graf21a, ), mean squared error (MSE) (han2022neural, ; tirer2022extendedunconstrainedfeaturesmodel, ; zhou2022optimizationlandscapeneuralcollapse, ), and supervised contrastive loss (pmlr-v139-graf21a, ). Combined with Prop. 1, these properties imply that label recovery via linear probing is guaranteed with few labels.
@@ 2 Related Work
Theoretical analyses of CL.  A growing body of work aims to explain the effectiveness of CL. Early studies attributed CL’s success to maximizing mutual information between augmentations of the same image (NEURIPS2019_ddf35421, ), but later work showed that enforcing strict bounds on mutual information can hurt performance (pmlr-v108-mcallester20a, ; Tschannen2020On, ). Another influential line of work focuses on the alignment and uniformity of learned representations (wang2020understanding, ; wang2021understanding, ; chen2021intriguing, ). For example, (wang2020understanding, ) showed that same-sample augmentations are aligned, while negatives are uniformly distributed on a sphere. chen2021intriguing extended this to a wider family of losses, showing how alignment and uniformity can be modulated by a hyperparameter.
While these are important properties, they do not characterize how CL algorithms organize samples from different classes in the embedding space. As a result, several papers (arora2019theoreticalanalysiscontrastiveunsupervised, ; tosh2021contrastive, ; zimmermann2021contrastive, ; pmlr-v151-ash22a, ; NEURIPS2021_2dace78f, ; NEURIPS2021_27debb43, ; 10.5555/3600270.3602219, ; pmlr-v162-shen22d, ; wang2022chaos, ; pmlr-v162-awasthi22b, ; pmlr-v162-bao22e, ) studied CL’s ability to recover meaningful clusters and latent variables in the data. However, these results often rely on strong assumptions about the data and augmentations, such as assuming augmentations of the same sample are conditionally independent given their cluster identity (see, e.g., (arora2019theoreticalanalysiscontrastiveunsupervised, ; pmlr-v162-saunshi22a, ; tosh2021contrastive, ; pmlr-v162-awasthi22b, )). In order to avoid such assumptions, haochen2023a studies function classes that induce a similar bias towards preserving clusters without any assumption on the connectedness of augmentation sets. However, their analysis is focused on minimizing a spectral contrastive loss which serves as a surrogate for the more practically used InfoNCE loss. pmlr-v195-parulekar23a instead proves that InfoNCE learns cluster-preserving embeddings by capping the representation class’s capacity so that any function that tries to carve a semantic cluster must also split an image from its own augmentation set, making cluster integrity the loss-minimizing choice.
@@
Other theoretical studies have analyzed SSL from different angles, such as feature learning dynamics in linear models and two-layer networks (NEURIPS2022_7b5c9cc0, ; 10.5555/3648699.3649029, ; pmlr-v139-wen21c, ; tian2023understanding, ), the role of augmentations (NEURIPS2020_4c2e5eaa, ; feigin2025theoreticalcharacterizationoptimaldata, ), the projection head (gupta2022understandingimprovingroleprojection, ; gui2023unravelingprojectionheadscontrastive, ; xue2024investigating, ; ouyang2025projection, ), CL’s sample complexity (alon2024optimal, ) and how to relax the dependence on large mini-batches DBLP:conf/icml/YuanWQDZZY22 . Other work explores connections between self-supervised contrastive and non-contrastive learning (wei2021theoretical, ; balestriero2022contrastive, ; NEURIPS2021_02e656ad, ; garrido2023on, ; shwartzziv2023an, ).
Neural collapse.  Neural collapse (NC) (doi:10.1073/pnas.2015509117, ; han2022neural, ) is a phenomenon that occurs in the final stages in training of overparameterized networks for classification. In this regime, we observe: (i) within-class collapse, where the embeddings of samples within each class converge to a single vector; (ii) class separation, where these class means spread out and often form a Simplex Equiangular Tight Frame (ETF); and a form of (iii) alignment between the feature space and the classifier’s weights.
To understand the emergence of neural collapse, several papers studied the emergence of these properties in different learning settings. In (mixon2020neuralcollapseunconstrainedfeatures, ; doi:10.1073/pnas.2103091118, ) they introduced the “unconstrained features model” or “layer peeled model”—where each sample’s embedding is treated as a free parameter in $\mathbb{R}^{d}$. Under such relaxations, research has demonstrated that neural collapse is a characteristic of global minimizers for many commonly used supervised loss functions (NEURIPS2022_cdce17de, ), including cross-entropy (LU2022224, ; 10.5555/3692070.3692468, ; pmlr-v139-graf21a, ), mean squared error (MSE) (han2022neural, ; tirer2022extendedunconstrainedfeaturesmodel, ; zhou2022optimizationlandscapeneuralcollapse, ), and supervised contrastive loss (pmlr-v139-graf21a, ; pmlr-v162-awasthi22b, ; gill2023engineeringneuralcollapsegeometry, ; kini2024symmetric, ). In Thm. 2 we use a similar approach in order to characterize the global minima of the CL and NSCL losses, connecting them to downstream performance using tools from (galanti2022on, ; galanti2022improved, ; galanti2023generalizationboundsfewshottransfer, ).
@@ 3 Theoretical Analysis
The above theorem gives a justification for why contrastive SSL yields class-aware features without ever seeing the labels. Interestingly, since the right-hand side depends only on the number of samples $N$ and the maximal number of samples per class $n_{\max}$—not on the specific assignment $Y$—the inequality is label-agnostic. Namely, for any labeling $Y$ with $n_{\max}$ samples per class, we have,
@@
where ($\mathcal{L}^{\mathrm{NSCL}}_{(Y)}(f)$ is the NSCL loss under that labeling. In particular, $C\to\infty$ the term $\log(1+\tfrac{n_{\max}\mathrm{e}^{2}}{N-n_{\max}})$ vanishes, the probability of drawing a same-class contrasting drops to zero, and DCL effectively collapses onto the NSCL for all labelings simultaneously.
@@ 3.2 Characterizing Good Representations for Few-Shot Learning
We formalize this via the expected $m$-shot classification error. Suppose we have a set of class-conditional distributions $D_{1},\dots,D_{C^{\prime}}$ and define $D=\frac{1}{C^{\prime}}\sum^{C^{\prime}}_{c=1}D_{c}$ as their uniform mixture. For example, $D_{1},\dots,D_{C^{\prime}}$ can represent a subset of $C^{\prime}\leq C$ classes from the pre-training task, where all $D_{i}$ are either training datasets or data distributions corresponding to their respective classes. Given an $m$-shot training set $\widehat{S}_{i}\sim D_{i}^{m}$ for each class $i$, we define:
$\text{err}_{m,D}(f)~{}:=~{}\mathbb{E}_{\widehat{S}_{1},\dots\widehat{S}_{C^{\prime}}}[\text{err}_{D}(h_{f,\widehat{S}})],$
where $h_{f,\widehat{S}}$ is a classifier trained using only the small labeled support $\widehat{S}=\cup^{C^{\prime}}_{c=1}\hat{S}_{c}$ and $\text{err}_{D}(h)$ is the probability that $h$ misclassifies a random test point drawn from the task distribution $D$. Specifically, we denote linear probing (LP) and nearest-class-center classification (NCCC) errors by $\textnormal{err}^{\textnormal{LP}}_{m,D}(f)$ and $\textnormal{err}^{\textnormal{NCC}}_{m,D}(f)$ which refer to $\text{err}_{m,D}(f)$ with $h_{f,\widehat{S}}$ being a linear classifier that minimizes $\textnormal{err}_{\widehat{S}}(h)$ and the nearest-class-center classifier $\arg\min_{c\in[C^{\prime}]}\|f(x)-\mu_{f}(\widehat{S}_{c})\|_{2}$.
Prior work (galanti2022on, ; galanti2022improved, ; galanti2023generalizationboundsfewshottransfer, ) shows that this error can be bounded in terms of how clustered the representations are. Specifically, for each class $i$, let $\mu_{i}=\mathbb{E}_{x\sim D_{i}}[f(x)]$ and $\sigma^{2}_{i}=\operatorname{Var}_{x\sim D_{i}}\bigl{[}\|f(x)-\mu_{i}\|^{2}_{2}\bigr{]}$ denote the mean and variance of the class embeddings, and let $V_{f}(D_{i},D_{j})=\sigma^{2}_{i}/\|\mu_{i}-\mu_{j}\|^{2}$ be the class-distance-normalized variance galanti2022on ; galanti2022improved ; galanti2023generalizationboundsfewshottransfer between classes $i$ and $j$.
@@
In supervised learning, representations tend to become tightly clustered, causing CDNV to decrease, and thus reducing few-shot error (doi:10.1073/pnas.2015509117, ; galanti2022on, ; galanti2023comparative, ; zhou2022are, ).
However, in the self-supervised case, labels are not available during training. As a result, there is no mechanism to explicitly encourage intra-class similarity, and thus we cannot generally expect that $\text{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]$ will be small.
To better capture settings where label information is implicit, we introduce a refinement of (3) that prioritizes directional variability. For each pair $i,j$, let: $\sigma_{ij}^{2}=\operatorname{Var}_{x\sim D_{i}}\bigl{[}\langle f(x)-\mu_{i},u_{ij}\rangle\bigr{]}$, where $u_{ij}=\frac{\mu_{i}-\mu_{j}}{\lVert\mu_{i}-\mu_{j}\rVert_{2}}$. Then the directional CDNV is defined as: $\tilde{V}_{f}(D_{i},D_{j})=\sigma_{ij}^{2}/\|\mu_{i}-\mu_{j}\|_{2}^{2}$, that measures the variation along the line between the class centers.
This leads to the following stronger bound:
@@
Let $C^{\prime}\geq 2$ and $m\geq 10$ be integers. Fix a feature map $f:\mathcal{X}\to\mathbb{R}^{d}$ and class-conditional distributions $D_{1},\dots,D_{C^{\prime}}$ over $\mathcal{X}$. We have:
$\mathrm{err}^{\mathrm{LP}}_{m,D}(f)~{}\leq~{}\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)~{}\leq~{}(C^{\prime}-1)\left[14\textnormal{Avg}_{i\neq j}[\tilde{V}_{f}(D_{i},D_{j})]+\tfrac{28.5}{m}\textnormal{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]\right].$ (4)
Unlike (3), which depends on the full CDNV, Prop. 1 depends primarily on the directional CDNV, (and on $\tfrac{1}{m}V_{f}$), which is always smaller than the CDNV ($\tilde{V}_{f}\leq V_{f}$) and in isotropic distributions $f\circ D_{i}$ it even scales as $\tilde{V}_{f}=\tfrac{1}{d}V_{f}$. Thus, the new bound can explain low $m$-shot error even when the regular CDNV is large. In essence, the bound predicts that SSL-trained models are effective for downstream tasks when their directional CDNV is very low and their CDNV remains moderate. Fig. 5 shows that this prediction holds in practice.

### v2

@@ 1 Introduction
Unsupervised representation learning refers to a class of algorithmic approaches designed to discover meaningful representations of complex data without relying on explicit supervision signals. The goal is to learn representations that preserve and expose semantic information, allowing them to be effectively leveraged in downstream supervised tasks. In recent years, these methods have proven to be effective in pre-training models on unlabeled data, enabling strong generalization on downstream computer vision tasks (hjelm2018learning; oord2019representationlearningcontrastivepredictive; NEURIPS2019_ddf35421; NEURIPS2020_70feb62b; pmlr-v119-chen20j; chen2020improvedbaselinesmomentumcontrastive; Chen_2021_ICCV; hénaff2020dataefficientimagerecognitioncontrastive; li2021prototypical; misra2019selfsupervisedlearningpretextinvariantrepresentations; 10.1007/978-3-030-58621-8_45; NEURIPS2020_4c2e5eaa; zbontar2021barlow; bardes2022vicreg; he2022masked; oquab2024dinov2learningrobustvisual) and in natural language processing (fang2020certcontrastiveselfsupervisedlearning; NEURIPS2020_1457c0d6; gao-etal-2021-simcse; su-etal-2022-tacl; reimers-gurevych-2019-sentence; zhang-etal-2021-pairwise; giorgi-etal-2021-declutr; yan-etal-2021-consert).
@@
Figure 1: DCL forms semantic clusters without label supervision, while NSCL yields tighter, more separable clusters, despite not explicitly pulling same-class samples together. We plot UMAP visualizations for (top) decoupled contrastive learning (DCL) (10.1007/978-3-031-19809-0_38) and (bottom) negatives-only supervised contrastive learning (NSCL) training on mini-ImageNet. See Appendix A for details.
One of the most successful paradigms for unsupervised learning is self-supervised contrastive learning (CL), where models are trained to distinguish different augmented views of the same image (positives) from views of other images (negatives), typically by minimizing the InfoNCE loss (pmlr-v9-gutmann10a; oord2019representationlearningcontrastivepredictive). Notable methods in this category include SimCLR (pmlr-v119-chen20j), MoCo (He_2020_CVPR; chen2020improvedbaselinesmomentumcontrastive; Chen_2021_ICCV), and CPC (oord2019representationlearningcontrastivepredictive). These approaches have led to representations that generalize remarkably well to downstream tasks, often competing with or even surpassing supervised learning. For example, an ImageNet-1M pre-trained MoCo achieves 81.5 $\mathrm{AP}_{50}$ on PASCAL VOC (journals/ijcv/EveringhamGWWZ10), while its supervised counterpart achieves 81.3 $\mathrm{AP}_{50}$.
Although the representations learned in supervised classification problems are relatively well understood–thanks to characterizations such as neural collapse doi:10.1073/pnas.2015509117; han2022neural, which occurs under a variety of loss functions including cross-entropy (LU2022224; 10.5555/3692070.3692468; pmlr-v139-graf21a), mean squared error (MSE) (han2022neural; tirer2022extendedunconstrainedfeaturesmodel; zhou2022optimizationlandscapeneuralcollapse), and supervised contrastive loss (pmlr-v139-graf21a; pmlr-v162-awasthi22b; gill2023engineeringneuralcollapsegeometry; kini2024symmetric)–the theoretical understanding of SSL remains limited. Surprisingly, despite the absence of labels, SSL-trained models often produce representations that closely resemble those learned via supervised training: they tend to cluster nicely with respect to semantic classes  (amir2021deep; shaul2023reverse; weng2025clusteringpropertiesselfsupervisedlearning) (see also Fig. 1) and support downstream tasks such as linear classification and clustering (pmlr-v119-chen20j; He_2020_CVPR; chen2020improvedbaselinesmomentumcontrastive; Chen_2021_ICCV; oord2019representationlearningcontrastivepredictive). This raises the following question:
@@
  • Duality between self-supervised and supervised CL: In Thm. 1 we formally establish a connection between decoupled contrastive learning (DCL) (10.1007/978-3-031-19809-0_38) and a supervised variant we call negatives-only supervised contrastive loss (NSCL), in which negative pairs are drawn from different classes. Our main insight is that, although DCL does not explicitly exclude same-class samples from the loss denominator, the probability of treating them as negatives vanishes as the number of classes increases. We prove that the gap between the two losses shrinks with more classes. Unlike prior analyses (arora2019theoreticalanalysiscontrastiveunsupervised; pmlr-v162-saunshi22a; tosh2021contrastive; pmlr-v162-awasthi22b; pmlr-v195-parulekar23a), our result holds without assumptions on model architecture, data distribution, or augmentations. Figs. 34 show that the two losses are highly correlated, their gap decreases as the number of classes increases, and minimizing the DCL loss implicitly minimizes the NSCL loss.
  • Estimating the few-shot error on downstream tasks: Prop. 1 introduces a bound on the $m$-shot classification error based on geometric properties of the learned representation. Specifically, the bound implies that lower class-distance-normalized-variance (CDNV) (see (galanti2022on) and Sec. 3.2) and directional CDNV (see Sec. 3.2) lead to lower $m$-shot error, with the impact of the regular CDNV diminishing as the number $m$ of per-class labeled examples increases. Empirically, we verify this behavior for DCL-trained models (Fig. 7) and find the bound to be tight in practice (Fig. 6).
  • Characterizing the global minima of the NSCL loss: In Thm. 2, we show that any global minimizer of the NSCL loss satisfies: (i) Augmentation collapse—all augmented views of a sample map to the same point; (ii) Within-class collapse—all samples from the same class share a representation; and (iii) Simplex ETF structure—class centers form a maximally separated, symmetric configuration. These solutions coincide with the optimal solutions of certain supervised classification settings, including those using cross-entropy loss (LU2022224; 10.5555/3692070.3692468; pmlr-v139-graf21a), mean squared error (MSE) (han2022neural; tirer2022extendedunconstrainedfeaturesmodel; zhou2022optimizationlandscapeneuralcollapse), and supervised contrastive loss (pmlr-v139-graf21a). Combined with Prop. 1, these properties imply that label recovery via linear probing is guaranteed with few labels.
@@ 2 Related Work
Theoretical analyses of CL.  A growing body of work aims to explain the effectiveness of CL. Early studies attributed CL’s success to maximizing mutual information between augmentations of the same image (NEURIPS2019_ddf35421), but later work showed that enforcing strict bounds on mutual information can hurt performance (pmlr-v108-mcallester20a; Tschannen2020On). Another influential line of work focuses on the alignment and uniformity of learned representations (wang2020understanding; wang2021understanding; chen2021intriguing). For example, (wang2020understanding) showed that same-sample augmentations are aligned, while negatives are uniformly distributed on a sphere. chen2021intriguing extended this to a wider family of losses, showing how alignment and uniformity can be modulated by a hyperparameter.
While these are important properties, they do not characterize how CL algorithms organize samples from different classes in the embedding space. As a result, several papers (arora2019theoreticalanalysiscontrastiveunsupervised; tosh2021contrastive; zimmermann2021contrastive; pmlr-v151-ash22a; NEURIPS2021_2dace78f; NEURIPS2021_27debb43; 10.5555/3600270.3602219; pmlr-v162-shen22d; wang2022chaos; pmlr-v162-awasthi22b; pmlr-v162-bao22e) studied CL’s ability to recover meaningful clusters and latent variables in the data. However, these results often rely on strong assumptions about the data and augmentations, such as assuming augmentations of the same sample are conditionally independent given their cluster identity (see, e.g., (arora2019theoreticalanalysiscontrastiveunsupervised; pmlr-v162-saunshi22a; tosh2021contrastive; pmlr-v162-awasthi22b)). In order to avoid such assumptions, haochen2023a studies function classes that induce a similar bias towards preserving clusters without any assumption on the connectedness of augmentation sets. However, their analysis is focused on minimizing a spectral contrastive loss which serves as a surrogate for the more practically used InfoNCE loss. pmlr-v195-parulekar23a instead proves that InfoNCE learns cluster‑preserving embeddings by capping the representation class’s capacity so that any function that tries to carve a semantic cluster must also split an image from its own augmentation set, making cluster integrity the loss‑minimizing choice.
@@
Other theoretical studies have analyzed SSL from different angles, such as feature learning dynamics in linear models and two-layer networks (NEURIPS2022_7b5c9cc0; 10.5555/3648699.3649029; pmlr-v139-wen21c; tian2023understanding), the role of augmentations (NEURIPS2020_4c2e5eaa; feigin2025theoreticalcharacterizationoptimaldata), the projection head (gupta2022understandingimprovingroleprojection; gui2023unravelingprojectionheadscontrastive; xue2024investigating; ouyang2025projection), CL’s sample complexity (alon2024optimal) and how to relax the dependence on large mini-batches DBLP:conf/icml/YuanWQDZZY22. Other work explores connections between self-supervised contrastive and non-contrastive learning (wei2021theoretical; balestriero2022contrastive; NEURIPS2021_02e656ad; garrido2023on; shwartzziv2023an).
Neural collapse.  Neural collapse (NC) (doi:10.1073/pnas.2015509117; han2022neural) is a phenomenon that occurs in the final stages in training of overparameterized networks for classification. In this regime, we observe: (i) within-class collapse, where the embeddings of samples within each class converge to a single vector; (ii) class separation, where these class means spread out and often form a Simplex Equiangular Tight Frame (ETF); and a form of (iii) alignment between the feature space and the classifier’s weights.
To understand the emergence of neural collapse, several papers studied the emergence of these properties in different learning settings. In (mixon2020neuralcollapseunconstrainedfeatures; doi:10.1073/pnas.2103091118) they introduced the “unconstrained features model” or “layer peeled model”—where each sample’s embedding is treated as a free parameter in $\mathbb{R}^{d}$. Under such relaxations, research has demonstrated that neural collapse is a characteristic of global minimizers for many commonly used supervised loss functions (NEURIPS2022_cdce17de), including cross-entropy (LU2022224; 10.5555/3692070.3692468; pmlr-v139-graf21a), mean squared error (MSE) (han2022neural; tirer2022extendedunconstrainedfeaturesmodel; zhou2022optimizationlandscapeneuralcollapse), and supervised contrastive loss (pmlr-v139-graf21a; pmlr-v162-awasthi22b; gill2023engineeringneuralcollapsegeometry; kini2024symmetric). In Thm. 2 we use a similar approach in order to characterize the global minima of the CL and NSCL losses, connecting them to downstream performance using tools from (galanti2022on; galanti2022improved; galanti2023generalizationboundsfewshottransfer).
@@ 3 Theoretical Analysis
The above theorem gives a justification for why contrastive SSL yields class-aware features without ever seeing the labels. Interestingly, since the right-hand side depends only on the number of samples $N$ and the maximal number of samples per class $n_{\max}$—not on the specific assignment YY—the inequality is label-agnostic. Namely, for any labeling $Y$ with $n_{\max}$ samples per class, we have,
@@
where $\mathcal{L}^{\mathrm{NSCL}}_{(Y)}(f)$ is the NSCL loss under that labeling. In particular, $C\to\infty$ the term $\log(1+\tfrac{n_{\max}\mathrm{e}^{2}}{N-n_{\max}})$ vanishes, the probability of drawing a same-class contrasting drops to zero, and DCL effectively collapses onto the NSCL for all labelings simultaneously.
@@ 3.2 Characterizing Good Representations for Few-Shot Learning
Refer to caption Refer to caption
(a) CDNV (b) dir-CDNV
Figure 2: Illustration of CDNV and directional CDNV. CDNV compares how tightly samples within a class cluster to how far apart class centers are; lower values indicate tighter clusters and larger gaps between classes. The latter measures variability only along the line connecting two class centers, highlighting the component most relevant for distinguishing those classes.
[…]
Prior work (galanti2022on; galanti2022improved; galanti2023generalizationboundsfewshottransfer) shows that this error can be bounded in terms of how clustered the representations are. Specifically, for each class $i$, let $\mu_{i}=\mathbb{E}_{x\sim D_{i}}[f(x)]$ and $\sigma^{2}_{i}=\mathbb{E}_{x\sim D_{i}}\bigl[\|f(x)-\mu_{i}\|^{2}_{2}\bigr]$ denote the mean and variance of the class embeddings, and let $V_{f}(D_{i},D_{j})=\sigma^{2}_{i}/\|\mu_{i}-\mu_{j}\|^{2}$ be the class-distance-normalized variance galanti2022on; galanti2022improved; galanti2023generalizationboundsfewshottransfer between classes $i$ and $j$.
@@
In supervised learning, representations tend to become tightly clustered, causing CDNV to decrease, and thus reducing few-shot error (doi:10.1073/pnas.2015509117; galanti2022on; galanti2023comparative; zhou2022are).
However, in the self-supervised case, labels are not available during training. As a result, there is no mechanism to explicitly encourage intra-class similarity, and thus we cannot generally expect that $V_{f}=\text{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]$ will be small.
To better capture settings where label information is implicit, we introduce a refinement of (3) that prioritizes directional variability (see Fig. 2 for visual illustration). For each pair $i,j$, let: $\sigma_{ij}^{2}=\operatorname{Var}_{x\sim D_{i}}\bigl[\langle f(x)-\mu_{i},u_{ij}\rangle\bigr]$, where $u_{ij}=\frac{\mu_{i}-\mu_{j}}{\lVert\mu_{i}-\mu_{j}\rVert_{2}}$. Then the directional CDNV, $\tilde{V}_{f}=\textnormal{Avg}_{i\neq j}[\tilde{V}_{f}(D_{i},D_{j})]$, is defined as: $\tilde{V}_{f}(D_{i},D_{j})=\sigma_{ij}^{2}/\|\mu_{i}-\mu_{j}\|_{2}^{2}$, that measures the variation along the line between the class centers. This leads to the following stronger bound:
@@
Let $C^{\prime}\geq 2$ and $m\geq 10$ be integers. Fix a feature map $f:\mathcal{X}\to\mathbb{R}^{d}$ and class-conditional distributions $D_{1},\dots,D_{C^{\prime}}$ over $\mathcal{X}$. Let $V^{s}_{f}=\textnormal{Avg}_{i\neq j}[V_{f}(D_{i},D_{j})]$. We have:
$\mathrm{err}^{\mathrm{LP}}_{m,D}(f)~\leq~\mathrm{err}^{\mathrm{NCC}}_{m,D}(f)~\leq~(C^{\prime}-1)\left[8\tilde{V}_{f}+\tfrac{8}{\sqrt{m}}V^{s}_{f}+\tfrac{8}{\sqrt{m}}V_{f}+\tfrac{4}{m}V_{f}\right].$ (4)
Unlike (3), which depends on the full CDNV, Prop. 1 depends primarily on the directional CDNV, (and on $\tfrac{1}{m}V_{f}$), which is always smaller than the CDNV ($\tilde{V}_{f}\leq V_{f}$) and in isotropic distributions $f\circ D_{i}$ it even scales as $\tilde{V}_{f}=\tfrac{1}{d}V_{f}$. Thus, the new bound can explain low mm‑shot error even when the regular CDNV is large. In essence, the bound predicts that SSL-trained models are effective for downstream tasks when their directional CDNV is very low and their CDNV remains moderate. Fig. 7 shows that this prediction holds in practice.